# Sharp Tools: How Developers Wield Agentic AI in Real Software Engineering Tasks

### v2

@@ IV-B2 Unsolicited Actions
We observed that the agent was sometimes excessively proactive in making changes to solve the issue, making changes beyond the prompt’s scope. We found that the agent made such unsolicited changes after 38% of participants’ prompts that did not seek code changes. In some cases, this was productive. For example, P5 mentioned “[the agent] anticipated what I wanted to do”. In other cases, this led to miscommunication. For example, when the agent made changes beyond what the prompt specified, P16 was forced to reject the change and write a follow-up prompt to specifically instruct the agent to only make the requested change. Another participant mentioned that the agent “always tries to generate new code, even though the right thing to do here might be [something else]”.
@@
We further observed that the agent ran terminal commands in 10% of prompts that did not request to run any. As further evidence that these commands were unwanted, participants were nearly three times more likely to stop the agent prematurely when its response included terminal commands (61% vs. 21%). Unlike in code changes, where the agent does not apply the changes until proposed and can undo changes, terminal executions can permanently change the state of the environment and the codebase. For example, while P12 was reviewing a test log file, the agent ran some new tests on the terminal, even though this was not instructed in the prompt. This led to the test log being deleted, derailing the participant’s work. Similarly, P11 had to abandon an issue due to the agent corrupting the code environment by running a long series of terminal commands.
IV-B3 Synchronicity
Parallel actions by the agent and the user sometimes led to challenges stemming from the participant’s unawareness of the actions being performed by the agent. While in theory simultaneous collaborations on the same codebase allowed for participants to never be blocked from directly working on the code, this did not always work out in practice. When P9 tried to work on code alongside the agent, it led to the agent duplicating a part of the fix, forcing the participant to reject the changes. To avoid such situations, some participants preemptively prevented themselves from working on the codebase. P10 expressed that “I want to make sure that the code changes are clear and I can see the diffs clearly so I’ll wait for it to stop thinking”. While the agent was iterating on a fix and seemed to be going down the wrong path, P11 noted similar concerns: “It is doing something which is no longer helpful… I am scared it will break something, and so I will wait for it to finish what its doing”. Pu and colleagues [34] similarly reported that proactive AI assistance can make users uncertain about what actions they can perform without disrupting the AI, and suggest addressing this by limiting the AI’s actions to those that do not interfere with the user’s immediate workflow.
IV-B4 Verbosity
Participants often sought and read explanations by the agent, doing so after 39% of the agent’s responses. However, several participants (P6, P7, P13, P15, P18) expressed frustration with the verbosity of such explanations. While P15 noted that “Most of the time when I’m coding, I don’t need detailed responses, only to see the changes made”, P16 mentioned of a response that “It’s giving me a lot of gibberish… I don’t need all this analysis”. Sometimes, verbosity even led to miscommunication. In the case of P15, due to the verbosity of the agent’s response, they did not notice that the response contained a documentation link that they were looking for, and instead searched the web manually for the link. Thus, while participants engaged with the agent’s explanations, these explanations did not always provide participants with the level of details they needed for effective communication. This observation echoes those made by Bansal and colleagues [32]
IV-B5 Sycophancy
@@
Sometimes, the tendency of the agent to agree with whatever the participant mentioned in their latest prompt led to a mistrust of the agent. In particular, when the agent produced a fix to an issue and an explanation for that fix, participants were hesitant to trust the agent when in subsequent responses it offered a seemingly contradictory explanation for the same issue. As P15 mentioned, “If I say that ’your change is wrong’, it will revert that change. That makes me lose confidence in it”. Similarly, P16 expressed in the post-study questionnaire that the agent “can be confused easily and mislead the developer so it can’t be trusted blindly”. Bansal and colleagues [32] similarly note that consistency in the agent’s outputs in critical to establishing trust in the agent. P18 applied an alternate approach that avoided such issues, explaining that “I tend to treat it like a human, where I don’t want to give it the answer, I want it to think about it, and come to a conclusion itself. It might have insights I don’t know myself”. They did this by asking the agent suggestive questions (e.g., are there any negative consequences of…) rather than instructing the agent. This allowed the agent to retrospect while making refinements rather than blindly changing its assumptions based on the participant’s latest instructions. This approach worked well for the participant - they were able to solve the issue as well as find additional related problems with the code (unnecessary type checking) that they did not know about previously.
IV-B6 Overconfidence
@@
Another aspect of the agent’s response that caused participants to mistrust it was the agent’s conviction that the changes it made in its latest response correctly solved the issue. As P9 noted in the post-study questionnaire, “It never stopped thinking it knew how to solve the problem even though it was wrong”. Even when the agent generated correct changes and explanations, its confidence in the alignments of these outputs to the issue sometimes led the participant to be hesitant of the agent. P18 noted about a generated fix that “it works, but its not really what I want”. Similarly, upon reviewing the agent’s description of files relevant to addressing the issue, P10 mentioned, “These things are true, but not necessarily things I want to modify. Khati and colleagues [48] similarly report that AI trust metrics for developers include not only correctness, but also alignment with their intentions.
@@
IV-B7 Ineffective Follow-Up Suggestions
The agent offered textual suggestions for follow-up prompts in 62% of its responses. These suggestions were usually not effective in advancing the conversation– participants replied to only 7% of such follow ups in their next prompt. One reason for this may be that these suggestions were often non-specific (e.g.: Figure 1d). When participants did reply, the suggestions were more specific (e.g.: Would you like me to show you how to remove the manual tracking code and let RichEdit handle double-clicks naively?). Further, while participants often needed refinements to code changes, the agent’s follow ups instead would often suggest moving forward to other steps in the issue resolution process. Thus, participants may also not have responded to follow ups because they did not always reflect the true state of the solution.
@@ IV-C RQ3: Factors Associated with Success
We considered participants to be successful in solving the issue if they were completely satisfied with the fix so as to submit it as a pull request (Complete Success, 10 issues) or if they were satisfied with the generated patch but needed to run some configuration steps or tests outside the scope of the study session before submitting the change as a pull request (Patch Success, 6 issues). Participants made progress but were unable to fully complete 10 issues (Incomplete), did not make any progress in 2 issues (No Progress), and abandoned 1 issue due to an inability to make any progress towards the fix (Abandon Fix). Although the total number of issues that participants worked on (N=33) is too low to conclusively reason about participants’ success in resolving issues, we observed that two main factors were associated with success.
@@ IV-C1 Collaboration Strategies
Participants were more likely to be successful in collaborations with the agent when they actively and regularly contributed to the problem-solving process.
For example, participants were more likely to successfully resolve issues when they communicated more with the agent – the average number of prompts used for successful issues was 10.3 as compared to 7.1 for unsuccessful issues. Refining the agent’s outputs was also a factor in success. Only in 19% of the successful issues did participants not request any code refinements from the agent. In addition, participants were only successful in 30% of the issues in which they did not ask for code refinements.
Notably, participants who applied the one-shot strategy (Section IV-A) were less successful in solving issues – they only succeeded in 38% of the issues for which they provided the agent with the entire issue description. On the other hand, participants who applied the incremental resolution strategy were successful in 83% of the issues they worked on. Similarly, participants who provided the agent with expert knowledge (Section IV-A2) tended to perform better than those who simply provided the agent with environmental context and thus left the entire task implementation to the agent without providing actionable feedback. Participants succeeded in 14 out of 22 (64%) issues in which they supplied such expert insights to the agent, as compared to only 2 out of the 7 (29%) issues in which they did not provide any expert knowledge.
Beyond providing insights to the agent, we also found that participants were more successful when they contributed to the code directly rather than just relying on the agent to generate the fix. Participants were successful in 79% (11 out of 14) of the issues in which they wrote some code manually, as compared to in 33% (5 out of 15) of the issues for which they did not write any code manually. This result did not extend to manual debugging and testing – participants who manually worked on debugging or testing code for an issue were successful in 53% (10 out of 19) of such issues, as compared to 60% (6 out of 10) for issues in which they did not perform manual debugging or testing.
@@ IV-C2 External Factors
Participants’ success was influenced by external factors relating to the tool and the task, such as prior experience and nature of the issue. The 3 participants with previous experience with Cursor had more success (3 out of 4 issues resolved), compared to others’ 52% (13 out of 25) success-rate.
Participants were most successful solving bugs that only concerned user interface changes (5 out of 5 issues resolved), followed by refactoring/variable renaming issues (2 out of 3 issues resolved), then feature requests (4 out of 8 issues resolved), and finally bugs (5 out of 13 issues resolved). These results suggest that participants may have struggled to collaborate with the agent on non-code generation tasks such as localization and debugging that are critical to bug resolution.
@@ V-A Collaborating Actively across the Development Process
A recent blog post by Anthropic [50] discusses the prevalence of the ‘feedback loop’ interaction pattern, in which users collaborate with agents to work on code tasks simply by providing the agent with feedback from the environment. In our study, we observed a similar interaction pattern, as many of participants’ prompts requesting code changes consisted only of environmental context or details of the original issue description. However, our results suggest that such interactions are less effective than those in which developers provide their own expert knowledge to the agent (Section IV-C1). Thus, when collaborating with agents, developers should not restrict their role to ‘human routers’ [51] that simply provide environmental feedback, but rather actively contribute to the problem-solving process to achieve better results. Since participants often took a managerial role in these collaborations (Section IV-A2), future work could explore how agents can adopt the communication strategies and behaviors exhibited by high-performing subordinate engineers.
Further, our results also suggest that developers should not restrict their active involvement to the parts of the software engineering process in which the agent is not particularly helpful, such as debugging and testing. Even though participants primarily used the agent to generate code changes, they were more successful when they collaborated with the agent by also manually writing code. On the other hand, while participants expressed misgivings about the agent’s ability to help with debugging (Section IV-A4), manually debugging and testing code did not appear to improve their likelihood of success.
Thus, developer-agent collaborations were most successful when both the developer and the agent actively contributed to issue resolution. In line with prior findings [33], as the agent was not as supportive in debugging and localizing, it was often unable to actively contribute to this part of the issue resolution process, and thus participants struggled to successfully resolve bugs. While software engineering benchmarks have clear, well set up tasks and environments for agents to work on, real development can involve messy environmental setup and debugging steps before the developer is ready to make code changes. However, agents often fail at these tasks, as reflected in their worse performance on more realistic benchmarks [4]. For SWE agents to be useful for developers in their day-to-day tasks, they must thus empower users across all the steps of the software engineering process, especially the messy and ambiguous ones.
@@ V-B Calibrating Output Scope
AI agents are designed to solve complex issues autonomously, and have the tools to do so [2]. However, we found that this design can backfire – in our study, participants who applied the one-shot delegation strategy were often unsuccessful (Section IV-C1). When the agent tries and fails to successfully implement complex changes, it is difficult for the developer to debug where the problems in the buggy patch arise from. Further, real-world complex issues often require expert knowledge that only an experienced developer might have. Yet, knowing what information could be relevant may be difficult for developers to foresee, and providing all such necessary information in a single prompt might be tedious for developers, particularly in the case of complex issues. Thus, interactive agents should leverage the presence of the user as a collaborator by internally calibrating the scope of their outputs when working on an issue. This could enable the developer to provide expert insights when needed, and effectively investigate any problems with the agent’s output. Further, such calibration enables the agent to have a clear idea of potential next steps and be aligned with the state of the solution, engendering actionable and relevant follow-up suggestions (Section IV-B7). Bajpai and colleagues [25] similarly found that in-IDE AI copilots are more effective when debugging if they focus on gathering information and communicating with the user before suggesting a fix. Such a pipeline, as also supported by [52], has the advantages of the incremental resolution strategy without requiring additional effort from the user to manually analyze the issue or intervene in the fix unless necessary.
@@ V-C Collaborating through Discussion, Not Sycophancy
Friction arising from the agent’s lack of tacit knowledge is difficult to anticipate for both the agent and the developer (Section IV-B1). This friction may occur regardless of how much context the agent gathers, which suggests that agents may always produce incorrect patches in some cases.
In contrast to pair programming, where developers spend considerable time discussing code logic and implementation details [53], the agent rarely engaged in problem-solving discussions. Instead, despite initial confidence in its changes, the agent immediately agreed with whatever the developer said in the latest prompt. While participants using incremental strategies were more successful, the agent contributes little to the solution quality in this paradigm since it simply follows participant instructions.
@@ VI Conclusion
In this paper, we’ve seen that developers can collaborate with a SWE agent to solve real-world software engineering tasks, but also that such a collaboration is fraught and that success is not guaranteed. While the capabilities of large language models and of the agents that build upon them expands at a dizzying rate, the capabilities of the humans who work with them are relatively static. The strategies, challenges, and communication patterns we document here are therefore likely to remain relevant even as SWE agents rapidly evolve. By understanding and addressing these enduring human factors, we hope this work contributes to a future where developer-agent collaboration is both effective and empowering.
A. Pre-Study Questionnaire
We asked the participants to answer the questions below before the study:
  1. 1.
    With which gender do you most identify?
    • Woman
    • Man
    • Non-binary / gender diverse
    • Prefer not to say
    • Other
  2. 2.
    What is your age?
    • 18-25
    • 26-35
    • 36-45
    • 46-55
    • 56-65
    • 66+
    • Prefer not to say
  3. 3.
    With which race/ethnicity do you most identify? (Select all that apply)
    • American Indian or Alaska Native
    • Asian
    • Central Asian
    • East Asian
    • South Asian
    • Southeast Asian
    • Black or African American
    • Native Hawaiian or Pacific Islander
    • White
    • Hispanic or Latino
    • Multiracial
    • Prefer not to say
    • Other
  4. 4.
    Which best describes your programming experience?
    Scale: Multiple Choice
    • 0 to 2 years professional programming experience
    • 3 to 5 years professional programming experience
    • 6 to 10 years professional programming experience
    • 11 to 15 years professional programming experience
    • More than 16 years professional programming experience
  5. 5.
    When coding, how frequently do you use the following tools to chat with AI?
    Scale: Never, Occasionally, Regularly
    • ChatGPT
    • GitHub Copilot Chat
    • Codeium Windsurf
    • Cursor
    • Tabnine
    • Claude
    (Scale for each: Never / Occasionally / Regularly)
  6. 6.
    If you have previously used any tools other than the ones mentioned above to chat with AI, please specify them below.
    Open-ended response
  7. 7.
    When coding, how frequently do you use the following tools’ AI-generated code completions?
    Scale: Never, Occasionally, Regularly
    • GitHub Copilot
    • Windsurf Cascade
    • Cursor AI
    • Tabnine
    • Blackbox AI
  8. 8.
    If you have previously used any AI tools other than the ones mentioned above for AI-generated code completions, please specify them below.
    Open-ended response
  9. 9.
    How frequently do you use AI developer tools?
    Scale: Never, Occasionally, Regularly
    • to write new code
    • to write documentation for existing code
    • to modify existing code
    • to explain unfamiliar code
    • to debug code
  10. 10.
    Please rate your agreement with the following statements. AI tools for fixing bugs:
    Scale: Not at all, Moderately, Extremely, Don’t Know
    • are easy to use
    • can understand your requests
    • can provide high quality responses
    • can help you to write better code
    • can help you to write code more quickly
    • are enjoyable to use
B. Post-Study Questionnaire
We asked the participants to answer the questions after the session with Cursor:
  1. 1.
    Please rate your agreement with the following statements. AI tools for fixing bugs:
    Scale: Not at all, Moderately, Extremely
    • are easy to use
    • can understand your requests
    • can provide high quality responses
    • can help you to write better code
    • can help you to write code more quickly
    • are enjoyable to use
  2. 2.
    To what extent did you view the AI tool as:
    Scale: Not at all, A little, Somewhat, A great deal
    • A reference guide
    • A content generator
    • A problem solver
    • A collaborator
    • A colleague
    • A coach
    • An advisor
    • An assistant
    • A reviewer
  3. 3.
    How important were these aspects of working with the AI tool?
    Scale: Not Important, Moderately Important, Very Important
    • Detailed textual explanations of code changes
    • Ability to see the steps that led the AI to the relevant files for the query
    • Ability to follow the steps the AI took as they happened in real time
    • Ability to reject and/or accept code changes proposed by the AI
    • Refining the outputs given by the AI by continuing the conversation
    • Ability to undo actions by resetting to the state of earlier queries in the chat
    • Ability of the AI to use the currently open file as context for its response
    • Ability to refer to files or code blocks for the AI to use as context
  4. 4.
    Did anything stand out to you about the experience of using the AI tool? For example, was anything good, bad, surprising, or notable?
    Open-ended response
C. Study Protocol
Following is the study protocol containing details of the tutorial task and guidelines for the tasks participants performed that the study administrator followed across all sessions:
Introduction
Hi, good morning/afternoon/evening. Thanks for taking time out to participate in this user study. Before we begin, could you fill a short pre-study survey? I’ll send over the link to the form, and you can fill it on your system. Your participant number is ____
Participant fills the pre-study questionnaire
Today, you will be using the AI in Cursor IDE by remotely controlling my system to work on issues in the ____ repository. You can choose to work on any issues that you feel you could solve by making direct edits to the codebase. The goal is for you to create a solution patch for the issue that you would feel confident attaching your name to and submitting as a pull request for review. If you get done solving the issue, we’ll choose another issue for you to work on. The session will be around an hour long, and you will fill out a short post-session survey at the end. As mentioned in the survey, I’ll be recording the meeting. Do you have any questions?
Great, let me share my screen.
Warm-up Task
Let’s go through a quick demo task so you are familiar with how the IDE and its AI features work. Cursor is a fork of VSCode, so it uses most of the shortcuts and functionalities present in VSCode. Today, we’ll be using the AI tool Cursor Agent. You can chat with Cursor Agent and ask it to implement code changes for you. Cursor Agent can read and analyse all any files in the IDE, make code edits, and also run terminal commands. Let’s try this out. I have cloned a GitHub repo (link) for a CLI-based python game that involves sorting balls into different containers. Let’s try out this game on the terminal by running python3 cli-ball-sort.py. Great, so as you can see, there are 5 different containers and 3 balls to sort. So now, let’s say we want to use Cursor Agent to modify the game. On the chat window on the right, I’ve entered a prompt (Prompt entered: change the number of containers to 6, the maximum number of balls in each container to 5, and test out the changes) that asks to change the number of containers to 6 and the number of balls to 5. Cursor uses the open file as context for its response, though it can also search through the codebase for other files as we’ll see. You can also manually add context to give to Cursor. For example, I can add some lines of this file (customizer.py) to the context. I can also remove context (remove those lines). Let’s try giving it our prompt.
Cursor Agent runs and makes edits
As you can see, the AI searched through the different files and made edits. You can now approve or reject these edits. Cursor has also run the game through the terminal, and it seems to be working. Let’s check how it did by running it ourselves.
Run the game on terminal
Great, since these changes look right, let’s approve them. Aside from rejecting these changes, you can also stop the agent at any point in its execution and restore to an earlier point in the conversation (demonstrate where this button is). Do you have any questions about Cursor?
Great, let’s get started with the study. After you’re doing solving an issue, I will save the git diff of the changes you made and revert the repository back to the original state. I’ll email you these changes to you after the session.
Main Task
The repository has been set up in Cursor along with the all the required installations on a new conda environment. As in the demo, you can use the terminal.
You are also free to access the internet. I have a browser window open with the open issues in this repository (open the browser window), you can use this window to access any resources you need.
Please try to think aloud as you go though your process. Try to use the AI for all your tasks in the study, but feel free to make small manual changes if necessary.
Just to reiterate, your objective in this session is to solve the issue efficiently to the best of your ability, not to evaluate or observe how well the AI can solve the issue.
Let me give you remote control of my screen. You can now look for an issue to start solving with Cursor Agent.
Reminders
Before moving on to a new issue, confirm if the participant is 100% happy with the change, and they can attach their name on the patch when submitting a PR. If not, ask them to keep working on the issue till they get to this point. Only move on in this case if it is not possible for them to reach this level of satisfaction with the issue in the scope of the study.
Choice of issue:
1st issue: Entirely participant’s choice
Subsequent issues: Work with participants to choose issues varying in difficulty and domain (e.g. feature vs bug) depending on how successful the first issue was.
Prompts to remind the participant to think aloud:
  • Can you explain what you’re trying to do right now?
  • Could you talk about what your thought process right now?
  • What are you thinking about/ trying to do?
  • Tell me a little what you’re doing
  • Could you explain how what you’re doing is relevant to the task?
Post-Task
Thanks for participating! The session was very insightful. I’ll send over the link to the post-study survey so you can fill it out on your system.
Participant fills the post-study questionnaire
Great! Before we end, do you have any questions?
Okay, then once again, thanks for participating! Have a nice day!

### v3

@@ IV-B2 Unsolicited Actions
We observed that the agent was sometimes excessively proactive, making changes beyond the scope of the participant’s prompt. We found that the agent made unsolicited code changes after 38% of participants’ prompts that did not seek any code changes. In some cases, this was productive. For example, P5 mentioned that the agent “anticipated what I wanted to do”. In other cases, this led to miscommunication. For example, when the agent made changes beyond what the prompt specified, P16 was forced to reject the change and write a follow-up prompt to specifically instruct the agent to only make the requested change. P7 mentioned that the agent “always tries to generate new code, even though the right thing to do here might be [something else]”.
@@
We further observed that the agent ran terminal commands in 10% of prompts that did not request to run any. As further evidence that these commands were unwanted, participants were nearly three times more likely to stop the agent prematurely when its response included terminal commands (61% vs. 21%). Unlike in code changes, where the agent does not apply proposed changes until explicitly accepted and can undo its changes, terminal executions can permanently change the state of the environment and the codebase. For example, while P12 was reviewing a test log file, the agent ran some new tests on the terminal, even though this was not instructed in the prompt. This led to the test log being deleted, derailing the participant’s workflow. Similarly, P11 had to abandon an issue due to the agent corrupting the code environment by running a long series of terminal commands.
IV-B3 Sychophancy
@@
Sometimes, the tendency of the agent to agree with whatever the participant mentioned in their latest prompt led to a mistrust of the agent. In particular, after the agent produced a diff and an explanation for that diff, participants were hesitant to trust the agent when in subsequent responses it offered a seemingly contradictory explanation for the same fix. As P15 mentioned, “If I say that ’your change is wrong’, it will revert that change. That makes me lose confidence in it”. Similarly, P16 expressed in the post-study questionnaire that the agent “can be confused easily and mislead the developer so [the agent] can’t be trusted blindly”. Bansal and colleagues [32] have also noted that consistency in the agent’s outputs in critical to establishing trust in the agent. P18 applied an alternate approach that avoided such issues, explaining that they “tend to treat [the agent] like a human, where I don’t want to give it the answer, I want it to think about it, and come to a conclusion itself. It might have insights I don’t know myself”. They did this by asking the agent suggestive questions (e.g., are there any negative consequences of…) rather than instructing the agent. This allowed the agent to retrospect while making refinements rather than blindly changing its assumptions based on the participant’s latest instructions. This approach worked well for the participant - they were able to solve the issue as well as find additional related problems with the codebase (unnecessary type checking) that they did not know about previously.
IV-B4 Overconfidence
@@
Kuttal and colleagues [40] report that AI agents for programming should be designed to respond with uncertainty when appropriate. Consistent with this finding, we found in our study that participants sometimes mistrusted the agent due to its certainty that the changes it made in its latest response correctly solved the issue. As P9 noted in the post-study questionnaire, “It never stopped thinking it knew how to solve the problem even though it was wrong”. Even when the agent generated correct changes and explanations, its confidence in the alignment of these outputs to the issue sometimes led the participant to be hesitant of the agent. P18 noted about a generated fix that “it works, but it’s not really what I want”. Similarly, upon reviewing a description of codebase files that agent suggested the user to analyze, P10 mentioned, “These things are true, but not necessarily things I want to modify [to solve the issue]”. Khati and colleagues [55] similarly report that AI trust metrics for developers include not only correctness, but also alignment with their intentions.
@@ IV-C RQ3: Factors Associated with Success
We considered participants to be successful in solving the issue if they were completely satisfied with the fix so as to submit it as a pull request (Complete Success, 10 issues) or if they were satisfied with the generated patch but needed to run some configuration steps or tests outside the scope of the study session before submitting the change as a pull request (Patch Success, 6 issues). Participants made progress but were unable to fully complete 10 issues (Incomplete), did not make any progress in 2 issues (No Progress), and abandoned 1 issue due to an inability to make any progress towards the fix (Abandon Fix). Although the total number of issues that participants worked on (N=33) is too low to conclusively reason about participants’ success in resolving issues, we observed two main factors associated with success. We summarize the effect of these factors in Table II, and describe them in detail below.
@@ IV-C1 Collaboration Strategies
Participants were more likely to be successful in collaborations with the agent when they took a more active role in this collaboration.
For example, participants were more likely to successfully resolve issues when they communicated more with the agent – the average number of prompts used for successful issues was 10.3 as compared to 7.1 for unsuccessful issues. Iterating on the agent’s outputs was also a factor in success – participants were only successful in 30% of the issues in which they did not ask for code refinements.
Notably, participants who applied the one-shot strategy (Section IV-A1) were less successful in solving issues – they only succeeded in 38% of the issues for which they provided the agent with the entire issue description. On the other hand, participants who applied the incremental resolution strategy were successful in 83% of the issues they worked on. Similarly, participants who provided the agent with expert insights (Section IV-A2) tended to perform better than those who simply provided the agent with environmental context. Participants succeeded in 14 out of 22 (64%) issues in which they supplied such expert insights to the agent, as compared to only 2 out of the 7 (29%) issues in which they did not provide any expert knowledge.
Beyond providing insights to the agent, we also found that participants were more successful when they manually wrote code rather than just relying on the agent to do so. Participants were successful in 73% of the issues in which they wrote code manually, as compared to in 36% of the issues for which they did not write any code manually. This result did not extend to manual debugging and testing – participants who manually worked on debugging or testing code for an issue were successful in 53% of such issues, as compared to 60% for issues in which they did not perform manual debugging or testing.
TABLE II: Participant Success vs Study Characteristics
Factor Attribute Success Rate
Collaboration Strategies
Delegation Strategy One-Shot 38%
Incremental Resolution 83%
Manual Actions Manual Code Writing/Editing 73%
No Manual Code Edits 36%
Insights Provided Expert Insights 64%
No Expert Insights 29%
Iteration Pattern Requested Code Refinements 68%
No Code Refinements Requested 30%
Conversation Length Less than 10 prompts 50%
10 prompts or more 64%
External Factors
Task Type Bugs 38%
Other Issues (Features, UI Fixes, etc.) 69%
Experience with Cursor Had Prior Experience 75%
No Prior Experience 52%
@@ IV-C2 External Factors
Participants’ success was influenced by external factors relating to the tool and the task, such as prior experience and nature of the issue. The 3 participants with previous experience with Cursor had more success (3 out of 4 issues resolved), compared to others’ 52% success-rate.
Participants were most successful solving bugs that only concerned user interface changes (5 out of 5 issues resolved), followed by refactoring/variable renaming issues (2 out of 3 issues resolved), then feature requests (4 out of 8 issues resolved), and finally bugs (5 out of 13 issues resolved). These results suggest that participants may have struggled to collaborate with the agent to work on steps specific to bug resolution such as localization and debugging.
@@ V-A Collaborating Actively across the Development Process
A recent blog post by Anthropic [57] discusses the prevalence of the ‘feedback loop’ interaction pattern, in which users collaborate with agents to work on code tasks simply by providing the agent with feedback from the environment. In our study, we observed a similar interaction pattern, as many of participants’ prompts requesting code changes consisted only of environmental context or details of the original issue description. However, our results suggest that such interactions are less effective than those in which developers provide their own insights to the agent (Section IV-C1). Thus, when collaborating with agents, developers should not restrict their role to ‘human routers’ [58] that simply provide environmental feedback, but rather actively collaborate with them to achieve better results. Participants in our study often took a managerial role in these collaborations, which suggests that agents may be more likely to perform the role of junior engineers in developer-agent software teams. This in turn suggests that the junior developer workforce may be more impacted by the introduction of AI into software development teams. Future work can explore how SWE agents can adopt the communication strategies and behaviors exhibited by high-performing subordinate engineers, and whether adopting these strategies affects developers’ perceptions of such agents.
Our results further suggest that developers should actively collaborate with the agent in all stages of the software engineering process, not only in those stages for which the agent is not particularly helpful, such as debugging and testing. Even though participants primarily used the agent to generate code changes, they were more successful when they collaborated with the agent by also manually writing code. On the other hand, even though participants expressed misgivings about the agent’s ability to help with debugging (Section IV-A4), manually debugging and testing code did not appear to improve their likelihood of success.
Thus, developer-agent collaborations were most successful when both the developer and the agent actively contributed to issue resolution. In line with prior findings [33], as the agent was not as supportive in debugging and localizing, it was often unable to actively contribute to this part of the issue resolution process, and thus participants struggled to successfully resolve bugs. While software engineering benchmarks have clear, well set up tasks and environments for agents to work on, real development can involve messy environmental setup and debugging steps before the developer is ready to make code changes. However, agents often fail at these tasks, as reflected in their worse performance on more realistic benchmarks [4]. For SWE agents to be useful for developers in their day-to-day tasks, they must empower users across all the steps of the software engineering process, especially the ones that are messy and ambiguous.
@@ V-B Calibrating Output Scope
AI agents are designed to solve complex issues autonomously, and have the tools to do so [2]. However, we found that this design can backfire – in our study, participants who applied the one-shot delegation strategy were often unsuccessful (Section IV-C1). When the agent tries and fails to successfully implement complex changes, it is difficult for the developer to debug where the problems in the buggy patch arise from. Further, real-world complex issues often require expert information that only an experienced developer might know. Yet, knowing what information could be relevant may be difficult for developers to foresee, and providing all such necessary information in a single prompt might be tedious for developers, particularly in the case of complex issues. Thus, interactive agents should leverage the presence of the developer as a collaborator by internally calibrating the scope of their outputs when working on an issue. This could enable the developer to provide expert insights when needed, and effectively investigate any problems with the agent’s output. Bajpai and colleagues [25] similarly found that in-IDE AI copilots are more effective when debugging if they focus on gathering information and communicating with the user before suggesting a fix. Such a pipeline, as also supported by [59], has the advantages of the incremental resolution strategy without requiring additional effort from the developer to manually analyze the issue or intervene in the fix unless necessary.
@@ V-C Collaborating through Discussion, Not Sycophancy
Friction arising from the agent’s lack of tacit knowledge is difficult to anticipate for both the agent and the developer (Section IV-B1). This friction may occur regardless of how much context the agent gathers, meaning that agents may always produce incorrect patches in some cases.
In contrast to pair programming, where developers spend considerable time discussing code logic and implementation details [60, 40], the agent rarely engaged in problem-solving discussions. Instead, despite initial confidence in its changes, the agent immediately agreed with whatever the developer said in the latest prompt. While participants using incremental strategies were more successful, the agent contributed little to the solution quality since it simply followed the user’s instructions.
@@ VI Conclusion
In this paper, we’ve seen that developers can collaborate with a SWE agent to solve real-world software engineering tasks, but also that such a collaboration is fraught and that success is not guaranteed. While the capabilities of large language models and of the agents that build upon them expands at a dizzying rate, the capabilities of the humans who work with them are relatively static. The strategies, challenges, and interaction patterns we document here are therefore likely to remain relevant even as SWE agents rapidly evolve. By understanding and addressing these enduring human factors, we hope this work contributes to a future where developer-agent collaboration is both effective and empowering.